# Loss Functions and Metrics in Deep Learning. A Review

### v1

@@ 4.2.6 F1-score
The F1 score considers both the model’s ability to correctly identify positive examples (precision) and the ability of the model to identify all positive examples in the dataset (recall). A higher F1 score indicates that the model has a better balance of precision and recall, whereas a low F1 score indicates that the model may have a high precision or recall but not both.
It is particularly useful when the class distribution is imbalanced, or we want to give equal weight to precision and recall.
4.2.7 F2-score
The F2 score is a variation of the F1 score, with more weight given to the recall metric. The F2 score is the harmonic mean of precision and recall, with a weighting factor of 2 for recall [68]. The formula for the F2 score is
$F2=(1+2^{2})\frac{Precision*Recall}{2^{2}*Precision+Recall}$ (30)
Like the F1 score, the F2 score ranges from 0 to 1, with a higher score indicating better performance. However, the F2 score places a greater emphasis on recall, making it useful when it is important to minimize false negatives. For example, a false negative could mean a patient is not diagnosed with a serious disease in medical diagnosis, so the F2 score is often used in such scenarios [69].
4.2.8 Specificity
@@
Specificity, also known as the true negative rate (TNR), is a metric that measures the proportion of actual negatives that are correctly identified as negatives by a classification model. It is defined as the number of true negatives (TN) divided by the number of true negatives plus the number of false positives (FP) [70]. The formula for specificity is
$Specificity=\frac{TN}{TN+FP}$ (31)
This metric is particularly useful in medical diagnostic testing, where it is important to minimize the number of false positives to avoid unnecessary treatments or interventions. High specificity indicates that the model is good at identifying negatives and has a low rate of false positives.
It is often used with the Recall or TPR to evaluate the overall performance of a classification model.
4.2.9 False Positive Rate (FPR)
The False Positive Rate (FPR) is used to evaluate the proportion of false positives (i.e., instances that are incorrectly classified as positive) to the total number of negatives (i.e., instances that are correctly classified as negative). It is also known as the Type I Error rate, which complements the Specificity metric.
Formally, the FPR is calculated as
$FPR=\frac{FP}{FP+TN}$ (32)
FPR directly relates to the threshold classifying instances as positive or negative. A lower threshold will increase the number of false positives and thus increase the FPR, while a higher threshold will decrease the number of false positives and decrease the FPR.
In practice, the FPR is often plotted on the x-axis of a Receiver Operating Characteristic (ROC) curve to visualize the trade-off between the TPR and FPR for different classification thresholds. See section 4.2.14 for more details.
4.2.10 Negative Predictive Value (NPV)
@@
The Negative Predictive Value (NPV) measures the proportion of negative cases that are correctly identified as such [70]. It is calculated as
@@
The NPV is useful when the cost of a false negative (i.e., an actual negative case being classified as positive) is high. For example, a false negative result in medical diagnostics can delay treatment or even death. In such cases, a high NPV is desired.
The NPV is not affected by the prevalence of the condition in the population, whereas other metrics, such as sensitivity and specificity, are. This makes the NPV a useful metric for evaluating the performance of a classifier when the class distribution is imbalanced.
The NPV can be interpreted as the complement of the false positive rate (FPR)
$NPV=1-FPR$ (34)
4.2.11 True Discovery Rate (TDR)
True Discovery Rate (TDR) evaluates the proportion of true positive predictions a model makes among all the positive predictions. It is also known as the Positive Predictive Value (PPV) or precision of the positive class [70]. TDR is calculated as
$TDR=\frac{TP}{TP+FP}$ (35)
TDR is a useful metric for evaluating the performance of a model in situations where the number of false positive predictions is high, and the number of true positive predictions is low. It is particularly useful in high-dimensional datasets where the number of features is large and the number of positive observations is low. TDR can provide a more accurate picture of the model’s performance than accuracy or recall in such cases.
There may be a trade-off between TDR and recall in some cases: TDR may be low when the recall is high, and vice versa. Therefore, it’s important to consider both TDR and recall when evaluating the performance of a model.
4.2.12 False Discovery Rate (FDR)
The False Discovery Rate (FDR) measures the proportion of false positives among all positive predictions made by a classifier [70]. It is defined as
@@ 4.2.12 False Discovery Rate (FDR)
The FDR can be an alternative to the False Positive Rate (FPR) when the cost of false positives and true negatives differs. It is particularly useful in cases where the number of false positives is more critical than the number of false negatives, such as in medical testing or fraud detection. A lower FDR value indicates that the classifier makes fewer false positive predictions.
4.2.13 Precision-Recall Curve
The precision-recall (PR) curve is a graphical representation of the trade-off between precision and recall for different threshold values of a classifier. Precision is the proportion of true positive predictions out of all positive predictions, while recall is the proportion of true positive predictions out of all actual positive instances. The precision-recall curve plots precision on the y-axis and recall on the x-axis for different threshold values of the classifier [68].
@@ 4.2.13 Precision-Recall Curve
    Start with a binary classifier that can predict a binary outcome and estimate the probability of the positive class. These probabilities are also known as scores.
@@
    For every possible threshold (from 0 to 1) on these scores, calculate the Precision (see Section 4.2.3) and the Recall (see Section 4.2.4).
@@
    Plot a curve with Recall on the X-axis and Precision on the Y-axis. Figure 2(a) shows an example of Precision-Recall curves for three models.
@@
Figure 2(a) shows the precision/recall curves for three models trained on the same data. The dashed line shows the ideal performance. Each model reports its Average Precision metric (see Section 5.2.1 for more details on Average Precision). In the following, we explain how to interpret PR curves.
@@
The closer the curve is to the top-right corner, the better the model’s performance. Ideal performance is indicated by a point at (1,1), which signifies perfect precision (no false positives) and recall (no false negatives). If the curve is closer to the top-right corner of the plot, it indicates that the model achieves a good balance of precision and recall for most threshold settings.
The area under the curve (AUC-PR) provides a single-number summary of the information in the curve. The maximum possible AUC is 1, which corresponds to a perfect classifier. A random classifier will have an AUC of 0.5. A model with a higher AUC is generally considered better.
Steepness of the curve. Ideally, we want the recall to increase quickly as precision decreases slightly, resulting in a steep curve. This steepness reflects a good balance between precision and recall. If the curve is less steep, we are losing a lot of precision for small increases in recall.
@@
Comparison of different models. We can compare the PR curves of different models to understand their performance. If the PR curve of one model is entirely above that of another, it indicates superior performance across all thresholds.
4.2.14 Area Under the Receiver Operating Characteristic curve (AUC-ROC)
@@
The Area Under the Receiver Operating Characteristic Curve (AUC-ROC) is a commonly used performance metric for evaluating the performance of binary classification models [68]. It measures the ability of the model to distinguish between positive and negative classes by plotting the true positive rate (TPR) against the false positive rate (FPR) at various threshold settings. The AUC-ROC is a value between 0 and 1, with 1 indicating a perfect classifier and a value of 0.5 indicating a classifier that performs no better than random guessing.
The Area under the Receiver Operating Characteristic curve (AUC-ROC) offers a single-value summary of the model’s performance across all possible threshold values. This measure is particularly valuable when comparing the performance of different models, as its assessment is independent of threshold choice.
However, in cases where the positive and negative class distributions are significantly imbalanced, the AUC-ROC, while still applicable, may not provide the most accurate performance representation. With a heavy imbalance, the ROC curve can appear overly optimistic, as a low false positive rate can still mean a large number of false positives if the total count of actual negatives is high, resulting in a misleadingly high AUC-ROC value.
In such imbalanced scenarios, the Precision-Recall (PR) curve and its corresponding area under the curve (AUC-PR) can often provide a more nuanced and accurate performance assessment. As PR curves focus more on detecting positive instances, often the minority class in an imbalanced dataset, they can deliver a more insightful evaluation of a model’s ability to detect positive instances, providing a more relevant representation of the model’s performance.
@@
    Start with a binary classifier that can predict a binary outcome and estimate the probability of the positive class. These probabilities are also known as scores.
@@
    For every possible threshold (from 0 to 1) on these scores, calculate the TPR (see Section 4.2.4) and the FPR (see Section 4.2.9).
@@
    Plot a curve with FPR on the X-axis and TPR on the Y-axis. Figure 2(b) shows an example of ROC curves for three models.
@@
Figure 2(b) shows the ROC curves for three models trained on the same data. The dashed line shows random performance. Each model reports its Area under the curve (AUC) in the legend. In the following, we explain how to interpret ROC curves.
TPR and FPR on each axis: The True Positive Rate (TPR) is used for the vertical axis. It measures the proportion of actual positives that are correctly identified as such. The False Positive Rate (FPR), also known as the fall-out or Probability of False Alarm, measures the proportion of actual negatives that are incorrectly identified as positives. The ROC curve plots the TPR vs. FPR at different classification thresholds. Lowering the classification threshold classifies more items as positive, thus increasing both False Positives and True Positives.
@@
Area Under the ROC Curve (AUC-ROC): AUC provides an aggregate performance measure across all possible classification thresholds. AUC-ROC of a model equals the probability that the model will rank a randomly chosen positive instance higher than a randomly chosen negative instance. Hence, the higher the AUC-ROC score, the better the model (from 0 to 1).
Diagonal line equals random guess: The diagonal line in the ROC curve plot has an AUC of 0.5 and represents a model with no discriminatory ability, i.e., one that predicts positives and negatives at random.
Towards the top-left corner: The more the curve sits in the top-left corner, the better the classifier, as it means the True Positive Rate is high and the False Positive Rate is low.
Compare Models: ROC curves are useful for comparing different models. The model with a higher AUC and its curve towards the top-left corner is generally considered better.
(a) Precision/Recall curve
(b) ROC Curve
@@
Figure 2: Precision/Recall and ROC Curves. (a) Shows the precision/recall curves for three models trained on the same data. The dashed line shows the ideal performance. Each model reports its Average Precision metric. (b) Shows the ROC curves for three models trained on the same data. The dashed line shows random performance. Each model reports its Area under the curve (AUC) in the legend
5 Object Detection

### v3

@@ 3.2.2 Classification Performance Metrics
This score considers both the model’s ability to correctly identify positive examples (precision) and its ability to identify all positive examples in the dataset (recall). A higher F1 score indicates a better balance between precision and recall, suggesting that the model is not overly skewed towards one metric at the expense of the other. This balance makes the F1 score especially useful in scenarios with imbalanced class distributions, where one class might dominate over others.
The main limitation of the F1 score is that it does not differentiate between the types of errors (false positives and false negatives beyond their proportional impact on precision and recall). Therefore, in certain contexts where the cost of false positives differs significantly from the cost of false negatives, other metrics may be more appropriate.
F-beta Score
An extension of the F1 score is the F-beta score, which introduces a flexibility in balancing precision and recall according to specific requirements of different applications. It is defined as:
$F_{\beta}=(1+\beta^{2})\cdot\frac{\text{Precision}\cdot\text{Recall}}{(\beta^{2}\cdot\text{Precision})+\text{Recall}}$ (41)
where $\beta$ is a positive real factor, placing more emphasis on recall when $\beta>1$ and more on precision when $\beta<1$. This customization allows the metric to be more closely tailored to the specific costs associated with false positives and false negatives in various use cases.
Specificity or True Negative Rate (TNR)
@@
This metric measures the proportion of negatives that are correctly identified as such by a classification model [65]. It is defined mathematically as the number of true negatives (TN) divided by the sum of true negatives and false positives (FP):
$\text{Specificity}=\frac{TN}{TN+FP}$ (42)
Specificity is used in fields like medical diagnostics, where a high value indicates that the test is effective at ruling out non-diseased individuals, thus minimizing unnecessary treatments or anxiety. This metric complements recall by focusing on different aspects of the testing outcome. Recall assesses the proportion of actual positives correctly identified, pertinent in scenarios where missing a positive case has severe implications, such as in cancer screening. Specificity, on the other hand, assesses the accuracy of identifying negatives, important in situations where falsely identifying a condition could lead to harmful interventions for example, in screening tests where a positive result could lead to more invasive testing or anxiety. It is also vital in industries like finance, where identifying a transaction as fraudulent when it is not can block legitimate transactions and potentially harm customer relations.
False Positive Rate (FPR)
The False Positive Rate (FPR) measures the proportion of false positives among all actual negative instances. It is also known as the Type I Error rate and is an important complement to the Specificity metric (also called True Negative Rate). The relationship between FPR and specificity is given by $FPR=1-\text{Specificity}$.
Formally, the FPR is calculated as follows:
$FPR=\frac{FP}{TN+FP}$ (43)
FPR is influenced by the threshold set for classifying instances as positive or negative. Lowering the threshold generally results in more instances being classified as positive, thus potentially increasing the FPR if many of these are actually negatives. Conversely, raising the threshold tends to decrease the FPR by classifying fewer instances as positive, reducing the risk of false positives.
In practice, FPR is commonly used in the construction of Receiver Operating Characteristic (ROC) curves, where it is plotted on the x-axis against the True Positive Rate (TPR) on the y-axis. This curve is useful for visualizing the trade-off between TPR and FPR at various threshold settings. The ROC curve and its analysis provide insights into the diagnostic ability of a classifier, regardless of the class distribution. See section 3.2.2 for more details on ROC curves and Area Under the Curve (AUC), which provides a single measure of overall model performance across multiple classification thresholds.
Negative Predictive Value (NPV)
@@
The Negative Predictive Value (NPV) measures the proportion of negative cases that are correctly identified as such. It specifically assesses the accuracy of negative predictions made by a classification model, calculated as follows:
@@
This metric is valuable where the consequences of false negatives are severe. For example, in medical diagnostics, a false negative might mean a delay in necessary treatment for a patient, potentially leading to worsened outcomes or even death. Thus, a high NPV is crucial in these situations.
Unlike some metrics such as sensitivity or specificity, NPV is directly affected by the prevalence of the condition being tested for, as it depends on the ratio of negatives (both true and false) in the dataset. This makes NPV especially useful in situations where negative outcomes (non-occurrence of the condition) are much more common than positive ones.
NPV complements the Positive Predictive Value (PPV) or precision, which measures the proportion of positive identifications that were actually correct. NPV and PPV together provide a comprehensive view of a model’s predictive accuracy regarding both positive and negative classifications.
In practice, NPV is useful in scenarios with imbalanced classes. For instance, in a disease screening where the disease is rare, the number of negatives (non-disease cases) will vastly outnumber the positives. In such cases, ensuring that the negatives are correctly identified (high NPV) without many false negatives becomes as important as identifying the positives.
The statistical relationship between the prevalence of a condition and NPV suggests that when the prevalence of a disease decreases, it becomes crucial to maintain a high NPV by controlling the number of false negatives. This is especially important in designing and evaluating screening tests for rare diseases where achieving a high NPV is often more meaningful.
False Discovery Rate (FDR)
The False Discovery Rate (FDR) measures the proportion of false positives among all positive predictions made by a classifier. This metric is crucial in fields where making incorrect positive predictions incurs significant costs or consequences. The formula for FDR is given by:
@@
A lower FDR value indicates that a smaller proportion of the positive predictions made by the classifier are false, signifying a higher reliability of positive results.
Unlike the False Positive Rate (FPR), which measures the proportion of false positives out of all actual negatives, the FDR focuses on the proportion of false positives within the group that the model predicts as positive. This distinction makes FDR particularly valuable in scenarios like genomic research or machine learning applications where the primary concern is the precision of the positive predictions rather than the impact on all negatives.
FDR is extensively used in scenarios where the cost of false positives is high, such as in scientific research where false discoveries can lead to incorrect scientific conclusions and waste of resources. In medical testing or fraud detection, controlling FDR ensures that the positive identifications are credible, which is crucial for subsequent decision-making processes.
Precision-Recall Curve
The precision-recall (PR) curve is a graphical representation that illustrates the trade-off between precision and recall for different threshold values of a classifier. Precision measures the proportion of true positive predictions out of all positive predictions, while recall measures the proportion of true positive predictions out of all actual positive instances. The PR curve plots precision on the y-axis and recall on the x-axis across different thresholds [64].
@@
    Begin with a binary classifier capable of predicting a binary outcome and estimating the probability of the positive class, often referred to as scores.
@@
    For every possible threshold (ranging from 0 to 1) applied to these scores, compute both Precision (see Section 3.2.2) and Recall (see Section 3.2.2).
@@
    Plot a curve with Recall on the X-axis and Precision on the Y-axis. An example is shown in Figure 2, depicting Precision-Recall curves for three different models.
@@
The effectiveness of a model as visualized by the PR curve can be interpreted as follows:
@@
  • Proximity to the top-right corner: The closer the curve approaches the top-right corner (1,1) of the plot, the better, indicating high precision and recall across most threshold settings. This ideal point represents a perfect classifier with no false positives or negatives.
  • Area under the curve (AUC-PR): The AUC provides a single-number summary of the model’s overall ability to achieve high precision and recall simultaneously. Unlike the ROC curve (see section 3.2.2), where a random classifier achieves an AUC of 0.5, the baseline AUC for a PR curve depends on the positive class prevalence.
  • Steepness of the curve: A steep initial rise of the curve is desirable as it indicates that the model achieves high recall with only a minimal reduction in precision, suggesting an effective balance between the two metrics.
@@
  • Model comparison: PR curves allow for the direct comparison of different models; a model whose PR curve lies consistently above another indicates superior performance at various threshold levels.
Figure 2: Precision/Recall curves for three models trained on the same data. The dashed line shows the ideal performance. Each model reports its Average Precision metric.
Precision-Recall Curve in Multi-Class Classification
While the Precision-Recall curve is traditionally used in binary classification settings, it can also be adapted for multi-class classification problems. In multi-class scenarios, a separate PR curve can be computed for each class, treating it as the positive class and all other classes combined as the negative class. This approach allows the evaluation of classifier performance for each specific class in a one-vs-all (OvA) fashion.
Computing Multi-Class PR Curves
To compute PR curves for multi-class classification, follow these steps:
  1. 1.
    For each class, treat the class as positive and all other classes as negative.
  2. 2.
    Calculate precision and recall for the class across a range of thresholds, just as in binary classification.
  3. 3.
    Plot a PR curve for each class, with recall on the x-axis and precision on the y-axis.
Each curve will provide insights into how well the classifier can identify instances of the respective class, while differentiating them from instances of all other classes.
Interpretation and Use of Multi-Class PR Curves
The interpretation of each PR curve follows the same principles as in binary classification:
  • Closer to the top-right corner: A curve that approaches the top-right corner indicates high precision and recall, showing effective class-specific performance.
  • Area under the curve (AUC-PR): The AUC for each class’s PR curve can be calculated to provide a summary statistic of the model’s performance concerning that class. A higher AUC indicates better overall performance in distinguishing the class from others.
Advantages and Limitations
This approach provides a detailed view of a classifier’s performance on a class-by-class basis, which is beneficial in settings where classes are imbalanced or have varying levels of importance. However, interpreting multiple PR curves can be complex, especially when classes are numerous or highly imbalanced. In such cases, additional summary statistics or visualization techniques may be necessary to provide a clearer overall performance picture.
To aggregate performance across multiple classes, macro-averaging and micro-averaging techniques can be used. Macro-averaging computes the average precision and recall across all classes, treating each class equally regardless of its prevalence. Micro-averaging aggregates the contributions of all classes to compute overall precision and recall, which can be particularly relevant in datasets where some classes dominate others.
Area Under the Receiver Operating Characteristic Curve (AUC-ROC)
@@
The Area Under the Receiver Operating Characteristic Curve (AUC-ROC) is a widely used metric for evaluating the performance of binary classification models. It measures the model’s ability to distinguish between the positive and negative classes across various threshold settings. The ROC curve plots the True Positive Rate (TPR) against the False Positive Rate (FPR) for different thresholds:
$\text{TPR}=\frac{TP}{TP+FN},\quad\text{FPR}=\frac{FP}{TN+FP}$ (46)
The AUC-ROC value ranges from 0 to 1, where 1 indicates perfect classification and 0.5 signifies no discriminative ability, equivalent to random guessing.
@@
    Begin with a binary classifier that predicts outcomes and estimates the probability of the positive class (referred to as scores).
@@
    Calculate the TPR and FPR at each threshold value from 0 to 1 (see Sections 3.2.2 and 3.2.2).
@@
    Plot these values with FPR on the X-axis and TPR on the Y-axis to form the ROC curve. An example of such a plot is shown in Figure 3.
@@
The ROC curve provides a graphical representation of a classifier’s performance. Key aspects of the ROC curve interpretation include:
  • Diagonal Line: Represents a random classifier. An ROC curve that lies above the diagonal indicates better than random performance.
@@
  • Area Under the Curve (AUC-ROC): Provides a single-value summary of the model’s effectiveness, indicating the probability that the classifier will rank a randomly chosen positive instance higher than a randomly chosen negative instance. Higher AUC values indicate better performance.
  • Toward the Top-Left Corner: The ideal position for an ROC curve, indicating both high TPR and low FPR.
  • Comparing Models: Models can be compared directly by their AUC values; a model with a higher AUC is generally better.
Limitations in Imbalanced Datasets
While AUC-ROC is a powerful metric, it may not fully represent performance in cases with significant class imbalance. The metric can appear optimistic as it primarily reflects the ranking of positive and negative instances without considering class distribution. In such scenarios, the Precision-Recall curve and its AUC (AUC-PR) might offer a more accurate performance assessment, especially focusing on the minority class.
@@
Figure 3: ROC curves for three models trained on the same data. The dashed line shows random performance. Each model reports its Area under the curve (AUC) in the legend.
Multi-Class Receiver Operating Characteristic Curve (AUC-ROC)
While ROC curves and the corresponding Area Under the Curve (AUC-ROC) are inherently designed for binary classification, they can be extended to multi-class problems through several approaches. These adaptations allow the ROC analysis to maintain its usefulness in evaluating classifiers that need to distinguish among more than two classes.
One-vs-All (OvA) Approach
The most straightforward method for extending ROC analysis to multi-class settings is the One-vs-All approach. Here, a ROC curve is constructed for each class by treating one class as the positive class and combining all other classes into a single negative class. This results in as many ROC curves as there are classes.
  1. 1.
    For each class, compute the True Positive Rate (TPR) and False Positive Rate (FPR) across various thresholds, treating that class as positive and all others as negative.
  2. 2.
    Plot these rates on the ROC curve, with FPR on the X-axis and TPR on the Y-axis for each class separately.
  3. 3.
    The AUC for each class-specific ROC curve can then be calculated, providing a measure of how well the classifier can distinguish each class from the rest.
One-vs-One (OvO) Approach
An alternative is the One-vs-One approach, where a ROC curve is generated for every pair of classes. This method is typically used in support vector machines and involves $(N\times(N-1))/2$ ROC curves for $N$ classes, which can become computationally intensive with a large number of classes but provides a very detailed analysis.
Macro and Micro-Averaging
To synthesize the information across multiple ROC curves in multi-class scenarios, macro-averaging and micro-averaging techniques are employed:
  • Macro-Averaging: Compute the AUC separately for each class’s ROC curve and then average these AUC scores. This method treats each class equally, regardless of class imbalance.
  • Micro-Averaging: Aggregate the contributions of all classes to compute an overall TPR and FPR across all thresholds, then plot a single ROC curve from these values. This approach gives a performance measure that reflects the classifier’s ability to distinguish all classes, weighted by class frequency.
Visualization and Interpretation
Visualizing multi-class ROC curves can be challenging due to the multiple curves involved. Effective strategies include:
  • Plotting each class’s ROC curve using different colors or line styles and providing a legend to help identify them.
  • Using panels or a grid layout where each panel represents the ROC curve for one class versus all others, or one pair of classes in the case of the OvO approach.
  • Providing summary statistics such as the mean AUC or the range of AUCs across classes to quickly gauge overall classifier performance.
Interpreting these curves requires considering the context of class importance and balance, as some classes may be more critical to classify correctly than others, or some classes may have more training examples, influencing their representation in the classifier’s training process.
4 Computer Vision Tasks
Up until now, we have covered loss functions and metrics for two fundamental tasks in Machine Learning, namely regression and classification. In the upcoming sections, we will go deeper into computer vision tasks such as image classification (section 4.1), object detection (section 4.2), image segmentation (section 4.3), face recognition (section 4.4), and generative models (section 4.5).
4.1 Image Classification
Image classification involves categorizing an image as a whole into a specific label [66].
Image classification was the first task where deep learning demonstrated significant success, marking a critical moment in the advancement of artificial intelligence. This breakthrough occurred primarily with the development and application of Convolutional Neural Networks (CNNs). In 2012, the AlexNet architecture [67], presented by Alex Krizhevsky, Ilya Sutskever, and Geoffrey Hinton, achieved state-of-the-art performance substantially better than the previous best results. This achievement established the potential of deep learning models to outperform traditional machine learning methods that relied heavily on hand-engineered features.
AlexNet’s success led to a surge in research and development in deep learning. This resulted in significant improvements in deep learning algorithms and their applications in various domains. Image classification played a key role in this revolution, enabling advances that expanded the scope of deep learning into numerous other applications. This breakthrough has laid a strong foundation for further innovations in the field of AI.
Image classification is used in a diverse range of sectors, demonstrating its broad applicability. In medical imaging, it is used to improve diagnostic accuracy by automatically detecting and classifying abnormalities such as tumors, fractures, and pathological lesions in various types of scans, including X-rays, MRIs, and CT images [68, 69, 70, 71]. In agriculture, it helps to monitor crop health, predict yields, and detect plant diseases through images captured by drones or satellites [72, 73, 74, 75]. The retail industry uses this technology for automatic product categorization, inventory management, and online shopping enhanced with visual search capabilities [76, 77, 78, 79]. Security and surveillance systems use image classification to analyze footage for unauthorized activities, manage crowds, and identify individuals or objects of interest [80, 81, 82]. In environmental monitoring, it is instrumental in tracking deforestation, monitoring changes in the natural environment, and assessing water quality using satellite imagery [82, 83, 84, 85, 86]. In addition, in the manufacturing sector, image classification automates quality control processes by identifying defects and verifying assembly, ensuring that products meet stringent standards with minimal human intervention [87, 88, 89, 90].
4.1.1 Image Classification Loss Functions
The loss functions used for image classification are the same as described in section 3.2.1, such as binary cross-entropy, categorical cross-entropy, weighted cross-entropy, focal loss, and hinge loss, as shown in Table 6.
4.1.2 Image Classification Metrics
The performance metrics used for image classification are the same described in section 3.2.2 and summarized in Table 7.
4.2 Object Detection